### v1

# Programming Quantum Measurements of Time inside a Complex Medium

Abstract
The temporal degree-of-freedom of light is incredibly powerful for modern quantum technologies, enabling large-scale quantum computing architectures and record key-rates in quantum key distribution. However, the generalized measurement of large and complex quantum superpositions of the time-of-arrival of a photon remains a unique experimental challenge. Conventional methods based on unbalanced Franson-type interferometers scale poorly with dimension, requiring multiple cascaded devices and active phase stabilization. In addition, these are limited by construction to a restricted set of phase-only superposition measurements. Here we show how the coupling of spatial and temporal information inside a single multi-mode fiber can be harnessed to program completely generalized measurements for high-dimensional superpositions of photonic time-bin. Using the multi-spectral transmission matrix of the fiber, we find special sets of spatial modes that experience distinct dispersive delays through the fiber. By exciting coherent superpositions of these spatial modes, we engineer the equivalent of large, unbalanced multi-mode interferometers inside the fiber and use them to perform high-quality measurements of arbitrary time-bin superpositions in up to dimension 11. The single fiber functions as a scalable, common-path interferometer for time-bin qudits that significantly eases the experimental overheads of standard approaches based on unbalanced Franson-type interferometers, serving as an essential tool for quantum technologies that harness the temporal properties of light.
I Introduction
The photonic temporal degree-of-freedom has recently emerged as a versatile property of light with powerful applications in quantum technologies. Photons traversing looped fiber architectures have enabled scalable approaches to universal linear optical quantum computing Rohde (2015), as well as powerful demonstrations of quantum computational advantage Madsen et al. (2022) and multi-photon interference Carosini et al. (2024). In parallel, photons encoded in discrete time-bins have been used to achieve record rates in quantum key distribution Islam et al. (2017) and noise-resilient entanglement distribution over large-scale distances Bulla et al. (2023); Yu et al. (2025). In this regard, the photonic time-domain acts as a high-dimensional quantum system that allows one to encode $d$-dimensional quantum states or qudits. Qudits have attracted significant interest in recent years as they provide a practical route towards high-capacity quantum networks Cozzolino et al. (2019); Mirhosseini et al. (2015) and device-independent quantum communication Srivastav et al. (2022); Vértesi et al. (2010). In addition, they provide powerful advantages for scalable quantum information processing by reducing circuit complexity and enhancing noise tolerance, with recent demonstrations in platforms ranging from photons Goel et al. (2024a); Lib et al. (2025), ions Meth et al. (2025), to superconducting circuits Champion et al. (2025).
The generalized measurement of large and complex superpositions of high-dimensional quantum states presents a unique experimental challenge. While measurement techniques for photonic qudits encoded in the transverse-spatial and path-integrated domains have seen significant development in recent years Goel et al. (2026); Forbes et al. (2025); Wang et al. (2020), scalable methods for measuring large time-bin qudits remain a major research endeavor. The archetypal time-bin qubit measurement utilizes a Franson-type unbalanced interferometer that coherently superposes an early and late time-bin to measure arbitrary time-bin qubit bases Franson (1989). Extending this method to $d>2$ dimensions is very difficult as it requires multiple cascaded and actively stabilized interferometers, as well as single-photon detectors with low timing jitter. As a result, time-bin measurements with Franson interferometers have been limited to $d=4$ with bulk optics Ikuta and Takesue (2017) and $d=8$ with integrated components in a recent experiment Yu et al. (2025). These limitations have inspired several alternative approaches for measuring high-dimensional time-bins that leverage random walks in coupled fiber loops Monika et al. (2025), diffraction in time via the temporal Talbot effect Widomski et al. (2024), and dispersion-engineered sum-frequency generation Serino et al. (2023). While promising, these techniques come with their own sets of limitations and technological challenges that have limited experimental demonstrations to dimensions $d\leqslant 5$.
Figure 1: $\tau$-modes of a multi-mode fiber (MMF): A random spatial mode $|r\rangle$ input into an MMF will generally scatter temporally and spectrally. The scattering behavior of an MMF is characterized by its multi-spectral transmission matrix $T(\omega)$, which can be Fourier-transformed to obtain its time-resolved transmission matrix $\tilde{T}(t)$. This matrix can be used to construct a special set of spatial $\tau$-modes ($|\tau_{0}\rangle$$|\tau_{1}\rangle$$|\tau_{2}\rangle$) that arrive at distinct times and are spectrally unscattered. A photon prepared in a coherent superposition of these spatial modes $|\tau_{X}\rangle=\frac{1}{\sqrt{3}}(|\tau_{0}\rangle+|\tau_{1}\rangle+|\tau_{2}\rangle)$ would arrive in a non-separable superposition of time and spatial modes $\frac{1}{\sqrt{3}}(|t_{0}\rangle|\tau_{0}\rangle+|t_{1}\rangle|\tau_{1}\rangle+|t_{2}\rangle|\tau_{2}\rangle)$.
Over the past decade, control over light scattering processes inside multi-mode fibers (MMFs) has led to a flurry of exciting advances in classical and quantum photonics Rotter and Gigan (2017); Plöschner et al. (2015). An MMF functions as a large, linear optical network for the photonic spatial, spectral/temporal, and polarization degrees-of-freedom (DoFs), and is characterized by its multi-spectral transmission matrix (MSTM) Andreoli et al. (2015); Mounaix et al. (2015). The ability to experimentally measure the MSTM of an MMF has enabled fundamental studies of principal, super-principal, and anti-principal modes—special states of light that experience strongly enhanced or reduced spectral correlations inside a multi-mode waveguide Carpenter et al. (2015); Ambichl et al. (2017). Alongside this, it has inspired interesting classical photonics applications such as spatial-temporal focusing of ultrashort pulses through an MMF Morales-Delgado et al. (2015); Mounaix and Carpenter (2019); Mounaix et al. (2020) and fiber-based spectrometers with ultrahigh spectral resolution Redding et al. (2014) . In the quantum realm, control over multi-mode light scattering in the spatial-polarization domains has been used to achieve programmable two-photon interference Defienne et al. (2016); Leedumrongwatthanakun et al. (2020), multi-mode entanglement transport Valencia et al. (2020), high-dimensional quantum circuits Goel et al. (2024b), and reconfigurable multi-user quantum networks Valencia et al. (2025) with a commercial MMF.
Figure 2: Programming generalized measurements of time-bin inside a multi-mode fiber (MMF). a) Schematic showing how a generalized, unbalanced Mach-Zehnder (Franson) interferometer can be used to measure an arbitrary 3-dimensional superposition of time-bins $|t_{in}\rangle$. The shapes (circle, triangle, square) are used as visual aids for labeling each time-bin. $|t_{in}\rangle$ traverses a color-coded, tunable phase-shifter $f_{i}$ and fixed delay in each arm, resulting in five peaks after the interferometer. The central peak contains a superposition of all three time-bins with desired phase-shifts introduced on each. In this manner, the Franson interferometer acts as a projective measurement of any phase-only time-bin superposition $|\psi\rangle=\frac{1}{\sqrt{3}}(f_{0}|t_{0}\rangle+f_{1}|t_{1}\rangle+f_{2}|t_{2}\rangle)$. Abbreviations: Single-mode fiber, SMF; Superconducting nanowire single-photon detector, SNSPD. b) In our alternative approach, a digital micromirror device (DMD) is used to shape the spatial mode of the time-bin superposition $|t_{in}\rangle$ into an arbitrary superposition of MMF $\tau$-modes depicted in the inset (green). As each $\tau$-mode experiences a different delay in the MMF, the fiber effectively functions as a three-mode unbalanced interferometer as in a), with path replaced by spatial mode. A spatial light modulator (SLM) after the MMF is used to coherently transform the output $\tau$-mode superposition into a Gaussian mode that is coupled into an SMF. The resulting five peaks are the same as those produced in a), except that the MMF approach does not require interferometric stabilization and allows for arbitrary (not just phase-only) superpositions to be measured. In addition, the number of time-bins and their separations can be tuned simply by using an MMF with a different core diameter and length, respectively (see text for more details).
Here we show how the coupling of spatial and temporal information inside an MMF can be harnessed to program generalized measurements for large photonic time-bin qudits. In particular, we use the fiber MSTM to construct a special set of spatial modes that arrive through the MMF unscattered in time. By exciting arbitrary coherent linear superpositions of these modes via wavefront shaping techniques, we are able to engineer large, multi-mode, unbalanced interferometers inside the MMF without the need for any phase stabilization. In this manner, the MMF functions as a generalized projective measurement in the time-bin basis, allowing us to perform time-bin measurements in mutually unbiased bases (MUBs) in dimensions up to $d=11$ inside a 40m-long commercial step-index MMF without any phase stabilizer. We certify the quality of our measurements by performing quantum tomography via a prepare-and-measure scheme that uses a tuneable Franson interferometer to prepare arbitrary time-bin qubits in all two-dimensional subspaces of our high-dimensional measurement Hilbert space. Our fiber-based method for realizing generalized, high-dimensional measurements of photonic time-bin can readily be scaled to larger time-bin dimensions while not requiring any active phase stabilization, making it particularly exciting for many quantum technology applications. Below, we present the central concept behind our technique, discuss its experimental implementation, and outline its advantages and current limitations.
II Concept
When a random spatial mode $|r\rangle$ localized at a specific time is input into an multi-mode fiber (MMF), the output mode is scattered both temporally and spectrally, i.e. the output spatial mode changes as a function of the input wavelength (see Fig. 1, top row). The scattering process in the spatial-spectral domains is captured by the fiber multi-spectral transmission matrix (MSTM) $T(\omega)$, which quantifies how a set of input spatial modes map to a set of output spatial modes at a given wavelength $\omega$. The MSTM can be Fourier-transformed to obtain the fiber’s time-resolved transmission matrix (TRTM) $\tilde{T}(t)$, which captures the spatial scattering behavior at a given time $t$. The TRTM can be used to construct a special set of orthogonal spatial modes that we call $\tau$-modes $|\tau_{i}\rangle$. The $\tau$-modes are obtained via a singular value decomposition (SVD) of the TRTM, and have been demonstrated experimentally for controllable energy deposition with ultrashort pulses Devaud et al. (2022). As seen in Fig. 1, the $\tau$-modes ($|\tau_{0}\rangle$$|\tau_{1}\rangle$$|\tau_{2}\rangle$) have the property that they arrive temporally and spectrally unscattered. However, the arrival time and spatial structure of each mode is different and is a function of the complex spatial-mode coupling and dispersion occurring within the MMF, characterized by the TRTM. A $\tau$-mode arriving at time $t$ is obtained from the singular vector corresponding to the largest singular value of the TRTM SVD at a given time $t$. The $\tau$-modes are equivalent to the set of principal modes under the assumption of first-order dispersion in the MMF (see Supplementary Section VI.4 for details on $\tau$-mode construction and comparison with principal modes). More importantly, due to linearity, the set of $\tau$-modes can be coherently combined to create new bases of modes. Fig. 1 depicts an example of such a superposition mode $|\tau_{X}\rangle=\frac{1}{\sqrt{3}}(|\tau_{0}\rangle+|\tau_{1}\rangle+|\tau_{2}\rangle)$ that would arrive in a non-separable superposition of temporal and spatial modes $\frac{1}{\sqrt{3}}(|t_{0}\rangle|\tau_{0}\rangle+|t_{1}\rangle|\tau_{1}\rangle+|t_{2}\rangle|\tau_{2}\rangle)$ after traveling through an MMF.
Figure 2a depicts a standard Franson interferometer modified for measuring arbitrary superpositions of three time-bins given by $|t_{\textrm{in}}\rangle=c_{0}|t_{0}\rangle+c_{1}|t_{1}\rangle+c_{2}|t_{2}\rangle$. The interferometer consists of three unbalanced paths with a fixed delay corresponding to the time difference $\Delta t$ between two adjacent time-bins. Each path additionally contains a tuneable phase-shifter $f_{i}$ that introduces a phase $e^{i\phi_{i}}$. After traversing the interferometer, the time-bin state $|t_{\textrm{in}}\rangle$ is split into three delayed copies, which are recombined to produce five time-delayed peaks. The central peak results from the interference of all three input time-bins, and its amplitude is proportional to the dot product of individual phase-shifter and time-bin probability amplitudes given by $f_{0}.c_{0}+f_{1}.c_{1}+f_{2}.c_{2}$. By setting the three phase-shifters to specific values $f_{i}$, the interferometer can be used to perform projective measurements of any phase-only time-bin superposition state given by $|\psi\rangle\langle\psi|$, where $|\psi\rangle=\frac{1}{\sqrt{3}}(f_{0}|t_{0}\rangle+f_{1}|t_{1}\rangle+f_{2}|t_{2}\rangle)$. Generalizing this device to $d$ dimensions requires $d$ unbalanced paths, resulting in $2d+1$ interference peaks. While here we have shown a multi-path interferometer for simplicity, in practice this is achieved by cascading multiple two-path interferometers Ikuta and Takesue (2017); Yu et al. (2025). Such implementations carry a significant overhead in terms of experimental complexity and stability, which has limited current demonstrations to $d\leqslant 8$. In addition, the cascaded approach does not give access to arbitrary $d$-dimensional superposition measurements.
Figure 2b depicts the central concept of our alternative approach. A time-bin superposition state $|t_{\textrm{in}}\rangle$ in a Gaussian spatial mode is incident on a digital micromirror device (DMD) that allows us to transform it into an arbitrary spatial mode before it is input into an MMF. We can choose this spatial mode to be an individual $\tau$-mode of the MMF, which would introduce a specific delay onto $|t_{\textrm{in}}\rangle$ (as shown in Fig. 1). By judiciously choosing this spatial mode to be in a coherent superposition $|\chi\rangle=\frac{1}{\sqrt{3}}(f_{0}|\tau_{0}\rangle+f_{1}|\tau_{1}\rangle+f_{2}|\tau_{2}\rangle)$ of three MMF $\tau$-modes, we can engineer the equivalent of a three-path unbalanced interferometer inside the MMF. A key difference from the bulk-optics approach in Fig. 2a is that here, the $\tau$-mode probability amplitudes $f_{i}$ can be complex, allowing us to implement tunable unbalanced interferometers that perform arbitrary (amplitude and phase) projective measurements of time-bin superposition states. Note that the delays introduced by the $\tau$-modes must be equivalent to the time-bin separations $\Delta t$, which ensures that the different copies interfere effectively. A spatial light modulator (SLM) at the output of the MMF is used to coherently transform the output $\tau$-mode superposition back into a fundamental Gaussian mode that is coupled into a single-mode fibre (SMF), leaving only the temporal interference information to be detected by a time-resolving detector (SNSPD).
The DMD-MMF-SLM setup is hence functionally equivalent to the multi-path Franson interferometer but utilizes the dispersive nature of the $\tau$-modes instead, giving us access to long delay times. The common-path interferometer implemented in this manner is also phase-stable and can readily be scaled to higher dimensions via an appropriate choice of fiber diameter and length. We experimentally realize this scheme with a 40-m-long commercial step-index MMF to implement programmable time-bin measurements in dimensions $d\in\{2,4,11\}$. First, we measure the MSTM of the MMF in the circularly polarized (CP) spatial-mode basis Plöschner et al. (2015); Snyder and Young (1978) by using a digital micromirror device (DMD) to generate 125 CP modes with a narrow-band swept laser source at $1550$ nmnm. These are launched into the MMF and the output complex field amplitude is reconstructed with an InGaAs camera using off-axis holography. We repeat this process for 421 wavelength steps of $3.8$ pmpm each, resulting in an MSTM with $421\times 125\times 125$ complex values. The measured MSTM is then Fourier-transformed to obtain a time-resolved transmission matrix (TRTM) with a resolution of 55 psps over a temporal range of $\approx 1.6$ nsns. Performing the SVD procedure on this TRTM results in 70 $\tau$-modes with a resolvable amplitude. From these, we choose 11 $\tau$-modes with a time-separation of $\approx 160$ psps to act as our time-bin basis (see Supplementary section V.1 for additional details).
III Experimental Results
We test the operation of our device as a programmable time-bin measurement by first using it to perform a tunable projective measurement of a time-bin qubit $|{t}_{L}\rangle=\frac{1}{\sqrt{2}}(|t_{0}\rangle-i|t_{1}\rangle)$ prepared with a standard, two-mode Franson interferometer. By setting the DMD/SLM to prepare/measure specific input/output $\tau$-mode superpositions in $d=2$, we program our device to perform a tunable, time-bin qubit measurement $|t(\theta)\rangle\langle t(\theta)|$, where $|t(\theta)\rangle=\frac{1}{\sqrt{2}}(|t_{0}\rangle+e^{i\theta}|t_{1}\rangle)$. In $d=2$, we expect our device to produce three peaks, with the counts contained in the central peak corresponding to the desired measurement. The resulting count rate as a function of relative phase $\theta$ is shown in Fig. 3. The visibility curve seen here depicts Malus’ law for polarization Malus (1809) applied to time-bin qubits. The minimum and maximum points on the curve indicate projective measurements of the orthogonal time-bin states $|{t}_{R,L}\rangle=\frac{1}{\sqrt{2}}(|t_{0}\rangle\pm i|t_{1}\rangle)$ with $\theta\in\{\pi/2,3\pi/2\}$, while the middle of the curve corresponds to measurements of the mutually unbiased basis states $|{t}_{D,A}\rangle=\frac{1}{\sqrt{2}}(|t_{0}\rangle\pm|t_{1}\rangle)$ with $\theta\in\{0,\pi\}$. The visibility of cosine-squared fit to the data in Fig. 3 is $0.97$, which shows the high extinction ratio of our time-bin “polarizer.”
Figure 3: Malus’ law for a time-bin qubit. A sinusoidal intensity pattern is observed when we prepare a time-bin qubit $|{t}_{L}\rangle=\frac{1}{\sqrt{2}}(|t_{0}\rangle-i|t_{1}\rangle)$ and implement a tunable measurement $|t(\theta)\rangle\langle t(\theta)|$ on it. The phase $\theta$ is varied by changing the $\tau$-mode superposition hologram on the DMD (see Fig. 2b). Sets of three interference peaks measured at the output when implementing time-bin measurements (i) $|t_{L}\rangle\langle t_{L}|$, (ii) $|t_{+}\rangle\langle t_{+}|$, and (iii) $|t_{R}\rangle\langle t_{R}|$ in our measurement device. The central peak corresponds to the probability of the measured time-bin state.
Next, we test the operation of our measurement device by implementing projective measurements of high-dimensional ($d\geqslant 2$) time-bin superposition bases known as mutually unbiased bases (MUBs). MUBs are of significant fundamental interest in quantum information science Tavakoli et al. (2021) and find many useful applications such as efficient entanglement witnesses and quantum state tomography Bavaresco et al. (2017); Guţă et al. (2020). In prime-power dimensions $d$, one can construct $d+1$ possible mutually unbiased bases (MUBs), with one possible construction given by Wootters and Fields (1989):
$|{v}^{\mu}_{a}\rangle$ $=\tfrac{1}{\sqrt{d}}\sum_{t=0}^{d-1}\omega^{at+\mu t^{2}}|t\rangle.$ (1)
Here, $a$ and $\mu$ label the state and basis respectively, $\omega=\exp(\tfrac{2\pi\hskip 0.5pti}{d})$ are the complex $d$-th roots of unity, and $|t\rangle$ are states in the time-bin (computational) basis. States from two different MUBs have an equal overlap of $1/d$ with each other.
Figure 4: Tomography of four-dimensional time-bin measurements. The left and right-most plots depict the phases and absolute values of 16 tomographically reconstructed measurement operators, $\hat{\mathcal{O}}$. These operators target the 16 possible MUB projectors for 4-dimensional time-bin superpositions $|{v}^{\mu}_{a}\rangle\langle{v}^{\mu}_{a}|$, which are shown in the middle for comparison. We observe an average fidelity of $96.24\%$ between the reconstructed measurement operators and their corresponding target projectors.
We perform projective measurements of these $d$-dimensional time-bin MUB states $|{v}^{\mu}_{a}\rangle$ by setting our device to display spatial MUBs of the corresponding input/output $\tau$-modes. Once the time-bin measurement is programmed, we characterize it using quantum tomography Hradil (1997); Qi et al. (2013); Smolin et al. (2012); Flammia et al. (2012). This involves probing the measurement device with a set of informationally complete time-bin states in $d=2$, and tomographically inverting the resulting data to obtain a density matrix that corresponds to the implemented measurement operator $\hat{\mathcal{O}}$. We can then calculate its fidelity to the target measurement $F=\langle{v}^{\mu}_{a}|\hat{\mathcal{O}}|{v}^{\mu}_{a}\rangle$. We use a tunable two-mode Franson interferometer to generate time-bin superpositions in every two-dimensional subspace of our $d$-dimensional space and input them into the measurement device. The Franson interferometer is phase-locked for stability and includes a $30$-cmcm-long motorized stage in one arm, allowing us to access time-delays across the full $1.6$ nsns range (see Supplementary sections VI.5 and V.2.4 for additional details on quantum tomography and the tunable Franson respectively).
We implement time-bin MUB measurements in dimensions $d\in\{2,4,11\}$. The reconstructed density matrices corresponding to all 16 time-bin MUB measurement operators in dimension $d=4$ are shown in Fig. 4. We don’t measure the computational (time-bin) basis as that is trivially obtained by using a time-resolving detector. The figure depicts the phase and magnitude of the reconstructed density matrix elements, alongside the target density matrices for comparison. As an example, the first row shows the four measurement operators corresponding to the target states in the first MUB $(\mu=1)$:
$|{v}^{1}_{1}\rangle$ $=(|t_{0}\rangle+|t_{1}\rangle+|t_{2}\rangle+|t_{3}\rangle)/2$
$|{v}^{1}_{2}\rangle$ $=(|t_{0}\rangle-|t_{1}\rangle+|t_{2}\rangle-|t_{3}\rangle)/2$
$|{v}^{1}_{3}\rangle$ $=(|t_{0}\rangle+|t_{1}\rangle-|t_{2}\rangle-|t_{3}\rangle)/2$
$|{v}^{1}_{4}\rangle$ $=(|t_{0}\rangle-|t_{1}\rangle-|t_{2}\rangle+|t_{3}\rangle)/2.$ (2)
Figure 5 (top row) shows reconstructed density matrices corresponding to 3 of the 132 possible time-bin MUB measurement operators in dimension $d=11$. These correspond to time-bin state superpositions in the first MUB, given by Eq. (1) with $\mu=1$ and $a\in\{1,2,6\}$. The target density matrices are shown in the bottom row. As can be seen in Figs. 4 and 5, the measured and target density matrices show good qualitative agreement with each other. To further quantify them, we calculate their fidelity to their target measurement operators. The average fidelities of the reconstructed time-bin measurement operators in each dimension is shown in Table 1. The measurement fidelity is seen to drop as dimension is increased. We attribute this to errors introduced in the tomography procedure, which involves the preparation of a large number of time-bin qubit probe states and probing the device over a long period of time. This introduces instabilities in the MSTM due to temperature fluctuations, which reduces the quality of the measurements as the dimension is increased.
Table 1: Average fidelities of reconstructed measurement operators in dimensions $d\in\{2,4,11\}$ to the ideal target measurements. The fidelity of each individual measurement can be found in the supplementary table 2. Errors are reported to one standard deviation.
Dimension 2 4 11
Fidelity 98.15 $\pm$ 0.39% 96.24 $\pm$ 0.30% 84.73 $\pm$ 1.27%
Figure 5: Tomography of eleven-dimensional time-bin measurements. The top row shows the angle and absolute values of three reconstructed time-bin MUB measurement operators $\hat{\mathcal{O}}$ in dimension $d=11$ ($|v^{1}_{1}\rangle\langle v^{1}_{1}|$, $|v^{1}_{2}\rangle\langle v^{1}_{2}|$ and $|v^{1}_{6}\rangle\langle v^{1}_{6}|$). Their corresponding target projectors, shown in the bottom row, are seen to agree qualitatively with the reconstructed measurement operators. We calculate a fidelity of $84.2\%$, $86.3\%$, and $83.7\%$ to the target measurements with a=1a=1,2, and 6 respectively.
IV Discussion and Outlook
We have proposed and demonstrated a method for measuring an arbitrary, high-dimensional time-bin superposition state by using the natural scattering processes inside a commercial multi-mode fiber (MMF). By precisely shaping an incoming time-bin superposition in its spatial structure, we can control its propagation in time through the MMF. This allows us to engineer the equivalent of large, multi-mode unbalanced interferometers inside the fiber, enabling its use as a programmable measurement device for arbitrary time-bin superposition states. We demonstrate the functionality of our device as a tunable projective measurement for mutually unbiased time-bin qudit states in up to dimension $d=11$ and reconstruct them via quantum tomography.
Our use of a single fiber is significant as it provides a simple pathway towards scaling such measurements to very large dimensions—a larger core diameter would support more spatial, and consequently, more time-bin modes and a longer fiber would ease timing jitter requirements. Above all else, the single fiber serves as a stable, common-path interferometer with large delays, surpassing the stringent alignment and stability requirements of bulk optical setups and the limited temporal delays possible on integrated platforms Ikuta and Takesue (2017); Yu et al. (2025). Interestingly, the standard approach based on cascading $N$ two-mode interferometers with $N$ independent phase controls for measuring $d=2^{N}$ dimensional states cannot access general transformations in dimension $d$. In contrast, our technique is completely general as it allows independent control over each time-bin amplitude.
Our technique currently suffers from a reduction in measurement quality as dimension is increased. While this appears to be a result of the temperature fluctuations over the time period required to tomographically reconstruct our measurement operators, it can be mitigated by incorporating analytic models for how small changes in temperature affect modal dispersion in the MMF Redding et al. (2014) and using an electro-optic modulator to prepare probe states in a fast and stable manner. In addition, the loss introduced during mode generation by the DMD due to its binary modulation can be reduced significantly by using a more efficient device such as a fast SLM or MEMS phase-only light modulator Rocha et al. (2024), as well as by incorporating more efficient mode conversion techniques Hiekkamäki et al. (2019). Our method serves as an exciting fundamental advance for quantum measurement science, where the measurement of large superpositions of time has been a long-standing challenge. It also holds significant potential for the measurement of high-dimensional time-bin and energy-time entanglement Xavier et al. (2025); Jha et al. (2008), making it particularly exciting for applications in long distance quantum networks with strong noise Srivastav et al. (2022); Bulla et al. (2023). Finally, it could find interesting applications in photonic quantum computing, where time-based architectures have recently attracted a lot of interest Bouchard et al. (2024); Humphreys et al. (2013); Takeda and Furusawa (2019).
Acknowledgements.
We would like to acknowledge David Phillips for helpful discussions and extend our thanks to Joel Carpenter for his library “digHolo” for performing high-speed off-axis holography. We acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC) (EP/Z533208/1), European Research Council (ERC) Starting Grant PIQUaNT (950402), and the Royal Academy of Engineering Chair in Emerging Technologies programme (CiET-2223-112).

### v2

@@
 In particular, we use the fiber MSTM to construct a special set of spatial modes that arrive through the MMF unscattered in time.
-By exciting arbitrary coherent linear superpositions of these modes via wavefront shaping techniques, we are able to engineer large, multi-mode, unbalanced interferometers inside the MMF without the need for any phase stabilization.
-In this manner, the MMF functions as a generalized projective measurement in the time-bin basis, allowing us to perform time-bin measurements in mutually unbiased bases (MUBs) in dimensions up to $d=11$ inside a 40m-long commercial step-index MMF without any phase stabilizer.
+By exciting arbitrary coherent linear superpositions of these modes via wavefront shaping techniques, we are able to engineer large, multi-mode, unbalanced interferometers inside the MMF.
+In this manner, the MMF functions as a generalized projective measurement in the time-bin basis, allowing us to perform time-bin measurements in mutually unbiased bases (MUBs) in dimensions up to $d=11$ inside a 40m-long commercial step-index MMF.
 We certify the quality of our measurements by performing quantum tomography via a prepare-and-measure scheme that uses a tuneable Franson interferometer to prepare arbitrary time-bin qubits in all two-dimensional subspaces of our high-dimensional measurement Hilbert space.
@@
 We would like to acknowledge David Phillips for helpful discussions and extend our thanks to Joel Carpenter for his library “digHolo” for performing high-speed off-axis holography.
-We acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC) (EP/Z533208/1), European Research Council (ERC) Starting Grant PIQUaNT (950402), and the Royal Academy of Engineering Chair in Emerging Technologies programme (CiET-2223-112).
+We acknowledge financial support from the UK Engineering and Physical Sciences Research Council (EPSRC) (EP/Z533208/1) (EP/Z533166/1), European Research Council (ERC) Starting Grant PIQUaNT (950402), and the Royal Academy of Engineering Chair in Emerging Technologies programme (CiET-2223-112).